# A Case Study on Misinformation Detection

### v3

@@ Acknowledgments
The authors would like to thank Anne Kohlbrenner for assisting with paper coding; Ben Kaiser and Sarah Scheffler for reviewing early drafts of this manuscript; and Mona Wang for retrieving Twitter data used for a replication study.
Availability
Our online SI is available in the following anonymous repo: https://anonymous.4open.science/r/sok_misinformation-B297/).
7 App: Full taxonomy of targets and critiques
{outline}
[enumerate] \1 Targets \2[ C] Claims \3[i.] Content-based detection via distance calculations on semantic embeddings \3[ii.] Content-based detection via search on knowledge graph topology \3[iii.] “Checkability” or “checkworthiness” \2[ A] Articles \3[i.] Syntactic and stylometric signals, including genre and sentiment \3[ii.] Topic-aware detection of stance and relevance to known rumoring topics \2[ U] Users \3[i.] Account metadata (bios, images, account age) \3[ii.] Single account behaviors (comments on posts, published posts) \2[ N] Networks \3[i.] Propagation patterns across social graphs \3[ii.] Timestamped records of user-user interactions \2[ W] Websites \3[i.] Text and URL/domain semantics \3[ii.] Site visitor demographics \3[iii.] Suspicious UI elements \3[iv.] Hosting infrastructure (DNS certificate, site age) \1 Datasets \2[i.] Dataset age (is it current or out of date?) \2[ii.] Leakage (evidence of temporal leakage, feature leakage) \2[iii.] Data dependencies (are they accounted/controlled for?) \2[iv.] Availability of information required to reproduce or reconstruct similar datasets (was non-public information required to produce ground-truth training sets?) \2[v.] Availability of original data \1 Models \2[i] Distance calculations on semantic embeddings \2[ii.] “Traditional” ML (SVM, RF, DT) \2[iii.] Deep learning (CNN, LSTM, GRU) \2[iv.] Graph cut algorithms \2[v] Stacked ensemble classifiers \2[vi.] Graph clustering algorithms \1 Features \2[i.] Textual \2[ii.] Network-based \2[iii.] Author-, user- or source-based \2[iv.] Infrastructural \1 Evaluation \2[i.] Testing in real time \2[ii.] Generalizability of approach \2[iii.] Evasion-resistance \2[iv.] Distributional shifts in training or test data \2[v.] Implications of high false-positive/false-negative rates
A Annotation Guide
Our guide for annotating papers follows below:
Target selection:
  • What is the stated target of detection?
  • What is the actual target of detection?
  • Is the detection method directly performing fact-checking and misinformation detection, or is it detecting proxy signals instead (e.g., emotional language, topics associated with rumoring narratives)?
Dataset curation:
  • Note dataset size, source, and temporality.
  • Note information scope: what is the “operative unit” of checkable information? Where this might become more complicated: some multimodal methods, particularly network-based ones, usually accept some combination of user- and post-scoped data, but only make a final decision about the veracity of one of those things.
  • Note information domain: does the data set coalesce around a topic / a set of topics? Are these pre-defined by the study authors?
  • Fine to name multiple scopes if one info scope is a vital parameter / input to the model in its determination of the veracity of another; e.g., a model that performs network analysis of social media posts that share a link to a specific news article; the model might make a final determination about the truthfulness of the news article, but propagation patterns + social media posts were a crucial element of that decision.
Feature set selection:
  • How were features selected?
  • Are features explainable and accessible? Were they handcrafted? Were they generated by unsupervised models?
  • Do positive signals indicate evidence of actually suspicious behavior?
  • Do authors account for feature noise?
Test-train split:
  • Note how/if this split was observed, and the percentage of the total data set assigned to test and training sets.
  • Note any overlap between both data sets (ideally, none).
  • Note any evidence of temporal leakage (future data used to make a prediction about past events).
Cross-validation performed:
  • Note size and number of folds.
  • Note presence of any development data sets (what percentage of total data set was this?).
  • Note presence of any hyperparameter tuning during cross-validation process – are these parameters published in-text?
Out-of-sample testing:
  • Note presence, if any.
  • What data sets were used for this?
  • Note any evidence of leakage during this process.
Third-party references:
  • Particularly for claim-based studies that require some form of ground-truth reference for fact-checking, what is the ground truth reference employed by the study?
  • Is this ground truth site reliable and up-to-date?
  • What kind of taxonomy does the ground-truth reference employ for labeling true / false statements?
  • What baked-in biases or slants might exist in the ground-truth reference’s data set (e.g., reference only labels political news sites; ground-truth reference is run by a known conservative organization)?
Model choice:
Evaluation:
  • Note dataset(s) used for evaluating classifier performance.
  • How does the evaluation dataset(s) compare to testing and training datasets in terms of size, sourcing, contents, and temporality?
  • How do authors account for distribution shifts in rumoring topics or activity types? Is there a road map for iterative updates to their method?

### v4

@@ Acknowledgments
The authors would like to thank Anne Kohlbrenner for assisting with paper coding; Ben Kaiser and Sarah Scheffler for productive conversations, and for reviewing multiple drafts of this manuscript; and Mona Wang for retrieving data used for a replication study.
[…]
Appendix
Appendix A Commercial fact-checking services
Media fact-checking. Human fact-checkers and content moderators affiliated with news outlets, or who work as freelance fact-checkers. The International Fact-Checking Network (IFCN) is a professional network of media workers and fact-checkers; IFCN is also the de facto standards setting body for media fact-checking, and maintains a fact-checking code of ethics [ifcn_code]. In general, media fact-checking organizations with IFCN affiliations are established news organizations, non-profits, and watchdog organizations that employ human journalists and fact-checkers. Furthermore, (human) fact-checkers can receive IFCN compliance certificates after passing a qualifying exam.
Brand safety and suitability companies. B2B companies that detect categories of potentially harmful speech on websites where ads might appear. Advertisers wishing to protect “brand safety” contract with these services to ensure that their ads do not appear alongside problematic content. The Global Alliance for Responsible Media (GARM) is the standards-setting body for brand safety and suitability companies [garm_standards].
  • Zefr, a GARM member company, deploys AI to detect material that falls within predefined subcategories of problematic content (e.g., explicit content, misinformation, spam). In a press release for Zefr’s acquisition of an AI-driven content moderation company (AdVerif.ai) from 2022, the company disclosed that AdVerif.ai is “powered by fact-checking data from more than 50 IFCN-certified organizations around the globe” [Zefr_2022]—that is, AdVerif.ai trains its models on labeled datasets produced by (human) IFCN affiliates.
  • DoubleVerify, a GARM member company, “uses sophisticated approaches that rely on a combination of AI and comprehensive human review” [Pasquine_2022]. According to the company’s documentation, human assessors (a “semantic science team”) evaluate site infrastructure and contents; AI is used to scale their assessments.
  • Integral Ad Science (IAS), a GARM member company, deploys AI to detect low-quality sites via infrastructure features. The company’s data sources, and deployment methodology were not immediately evident upon web search; IAS recently announced a new partnership with Meta for ad placement management on Facebook [Integral_Ad_Science_2024].
  • Twitter has deployed a crowd-sourced annotations platform called Community Notes (formerly Birdwatch) since 2021 [wojcik2022birdwatch].
  • Until recently, Facebook partnered with IFCN affiliates to perform third-party manual checking of possibly misinformative content; first-line automated methods detect potentially harmful speech and surface near-duplicates of known problematic image (SimSearchNet++) and text content [facebook_files_2021, fb_third_party, AI_at_Meta_2020]. Meta announced in January 2025 that it was sunsetting its third-party checking program in favor of a Community Notes-like system [chan_meta_2025].
  • TikTok employs thousands of content moderators across the globe who “work alongside automated moderation systems” [McIntyre_Bradbury_Perrigo_2023, Farah_2023].
  • Mandiant strongly implies the use of hybrid detection methods, and disclaims that “defenders must constantly explore different techniques and leverage both subject matter expertise and technical capabilities to filter and uncover malicious activity”) [Kapellmann_Zafra_Serabian_Riddell_Brubaker_2022].
  • Microsoft Threat Intelligence strongly implies the use of hybrid detection methods; in a report from September 2023, MTI cites the work of in-house “Microsoft Security teams” which are tracking an advanced social engineering attack [Microsoft_Incident_Response_2023]. Other details—including possible use of automated methods—are undisclosed.
  • Facebook Coordinated Inauthentic Behavior reports share quarterly updates about Meta’s takedown of coordinated activities across its platforms and others, including local news outlets. In a report from February 2023, Meta describes a CIB network in Serbia that used local news media to create the impression of grassroots support for the Serbian Progressive Party; while the nature of the detection methodology is unspecified, the complexity and geographic specificity of the CIB described suggest that specialists with country-level expertise were likely consulted [Nimmo_2023a].
  • The Global Disinformation Index (GDI) “reviews news domains based on various metadata and computational signals.” Content, however, is manually reviewed by a “country expert,” who analyzes a random sample of 10 articles from a news site to determine veracity [Srinivasan_GDI].
  • DFRLabs (Digital Forensic Research Lab) has disclosed that it employs human subject-area experts, and primarily addresses technology and policy issues pertaining to global and international affairs. In 2018, Facebook contracted its services to detect online trolls [Lapowsky_2018].
  • Graphika Labs leverages network analysis to identify influence operations online. On its own website and in the popular press, Graphika has disclosed that it uses AI to map online networks and trace information flows [Schiffrin_2023, graphika_ai].
LLM-driven detection. A few LLM-powered detection methods have been discussed in the popular press, including those advertised by Google [Storey_2023] and OpenAI [fried_2023], but these deployments appear to be mostly experimental, or have required additional adjudication from human moderators. OpenAI in particular has advertised content moderation tools that address misinformation-adjacent tasks, such as toxic speech detection [perspective_openai]. Misinformation and toxic speech detection are not equivalent tasks, however, and the latter is narrowly defined in the Perspective training data documentation as a four-way classification task (the four class labels are “profanity/obscenity,” “identity-based negativity,” “insults,” and “threatening” language).
Paper 1 Target 2 Dataset 3 Model 4 Features 5 Performance
Work Scope
Textual
Network
Author
Infra.
Accuracy/AUROC
1. Ajao et al. [ajao_sentiment_2019] C A Sentiment (A.i) PHEME [pheme_2018] LSTM, DT, RF, SVM 0.86 (Acc.)
2. Abulldah-All-Tanvir et al. [abdullah-all-tanvir_detecting_2019] C A N Content (C.i) Twitter (API) NB, RNN, LSTM, SVM, Logit 0.89 (Acc.)
3. Bhutani et al. [bhutani_fake_2019] C A Content (C.i); sentiment (A.i) Twitter (API), PolitiFact [politifact] Naive Bayes, RF 0.60 (AUC)
4. Bozarth et al. [bozarth-truth-2020] C Contents (C.i) PolitiFact [politifact], Daily Dot, Zimdars, MBFC LDA n/a
5. Ciampaglia et al. [ciampaglia-wiki-2015] C N Shortest path search (C.ii) DBpedia kNN, RF 0.97 (AUC)
6. Cui et al. [cui_same_2019] C A U Content (C.i); sentiment (A.i); user response (U.ii) PolitiFact[politifact], GossipCop[gossipcop] KNN, SVM, CSI [ruchansky_csi_2017], RMSprop 0.82 (F1)
7. Debnath et al. [debnath_hierarchical_2020] C Content (C.i) LIAR [wang_liar_2017] CNN 0.27 (Acc.)
8. Dey et al. [dey_fake_2018] C A Content (C.i); sentiment (A.i) Twitter (API) Clustering (kNN) 0.67 (Acc.)
9. Galitsky et al. [galitsky_detecting_2015] C Content (C.i) Amazon reviews Parse thicket 0.81 (Prec.)
10. Glockner et al. [glockner_missing_2022] C Content (C.i) PolitiFact[politifact], Snopes[snopes], MultiFC CNN, DNN 0.58 (Acc.)
11. Gordon et al. [gordon_disagreement_2021] C A Content (C.i); source rep (U.i) Credibility-Factors2020 SVD 0.63 (Acc.)
12. Gupta et al. [gupta_tweetcred_2015] C A N Content (C.i); stance (A.iii) Twitter (API) SVM 0.60 (Agreement)
13. Hassan et al. [hassan_data_2014] C Content (C.i); checkability (C.iii) NBA, weather datasets Frequency n/a
14. Jain et al. [jain_towards_2016] C Content (C.i); sentiment (A.i) Twitter (API) Gensim/TextBlob 0.77 (Acc.)
15. Jiang et al. [jiang_linguistic_2018] C Content (C.i); ling./syntax (C.ii) PolitiFact[politifact], Snopes[snopes] SVM 0.81 (Acc.)
16. Karimi et al. [karimi_multi-source_2018] C Content (C.i) LIAR[wang_liar_2017] LSTM, CNN 0.39 (Acc.)
17. Kartal et al. [kartal_too_2020] C Content (C.i); checkability (C.iii) Check That! dataset Logit, SVM, RF 0.26 (MAP)
18. Kou et al. [kou_hc-covid_2022] C Content (C.i) CoAID, CONSTRAINT Knowledge graph 0.90 (Acc.)
19. Paudel et al. [paudel_lambretta_2023] N A Keyword detection (A.ii) Abilov et al. dataset, Twitter (API) AdaRank, ListNet, RF 0.79 (MAP)
20. Popat et al. [popat_credibility_nodate] C Content (C.i) PolitiFact[politifact], Snopes[snopes], NewsTrust biLSTM, CNN 0.88 (AUC)
21. Shiralkar et al. [shiralkar_finding_2017] C N KG search (C.ii) DBpedia Knowledge graph 1.00 (AUC)
22. Shu et al. [shu_defend_2019] C U Word encoding (C.i); user response (U.ii) GossipCop [gossipcop], PolitiFact [politifact] RNN/RMSprop, CSI [ruchansky_csi_2017], LSTM, CNN 0.93 (F1)
23. Tian et al. [tian_early_2020] C U Content (C.i); user response (U.ii) Twitter15, Twitter16 CNN-biLSTM 0.82 (F1)
24. Zhang et al. [zhang_reply-aided_2019] C U Content (C.i); user response (U.ii) RumourEval, PHEME[pheme_2018] biLSTM, Multitask, SVM, CNN, 0.89 (Acc.)
1. Afroz et al. [afroz_detecting_2012] A Content (C.i); syntax (A.i) Brennan-Greenstadt SVM, J48 Decision Trees 0.97 (F1)
2. Ahmed et al. [ahmed_detecting_2018] A Syntax (A.i) Twitter, Kaggle, Horne and Adali [horne_this_2017] SVM 0.92 (Acc.)
3. Bourgonje et al. [bourgonje_clickbait_2017] A Stance (A.ii) Fake News Challenge Data Logit 0.90 (Acc.)
4. Brasoveanu et al. [brasoveanu_semantic_2019] A C Sentiment (A.i); keywords (A.ii) LIAR[wang_liar_2017] CNN, LSTM, CN 0.64 (Acc.)
5. Della Vedova et al. [della_vedova_automatic_2018] A N Content (C.i); virality (N.iv) FakeNewsNet, Buzzfeed Logit 0.82 (Acc.)
6. Horne et al. [horne_this_2017] A Syntax (A.i), headline (A.ii) Buzzfeed, Burfoot & Baldwin SVM 0.78 (Acc.)
7. Jabiyev et al. [jabiyev_fade_2021] A W Topic detection (A.ii); site cred. (W.iv) Snopes[snopes], FactCheck, PolitiFact[politifact] SVM, DT, RF 0.87 (Acc.)
8. Jadhav et al. [jadhav_fake_2019] A Content (C.i); syntax (A.i) LIAR [wang_liar_2017] DSSM/RNN 0.99 (Acc.)
9. Jin et al. [jin_detection_2017] A N U (A.ii); (N.i); suspicious accounts (U.i) Tweets; articles n/a 0.87 (Prec.)
10. Kapusta et al. [kapusta_fake_2020] A Sentiment & word freq. (A.i) MBFC and custom n/a n/a
11. Kumar et al. [kumar_disinformation_2016] A W U (A.i); UI (W.iii); Author cred. (U.i) 20K Wiki Hoaxes Random forest 0.87 (AUC)
12. Magdy et al. [magdy-stat-2010] A Content (C.i) NYT Corpus [nyt_corpus], 100 Wikis Pattern recog. 0.99 (Recall)
13. Monti et al. [monti2019fake] A N U (A.ii); (N.i); suspicious accounts (U.i) Tweets; articles RNN/CNN 0.927 (AUC)
14. Nasir et al. [nasir_fake_2021] A Syntax (A.i) ISOT [isot]; FAKES [fa_kes] RNN/CNN 0.99 (Acc.)
15. Perez-Rosas et al. [perez-rosas_automatic_2017] A Syntax (A.i) FakeNewsAMT; Celebrity SVM 0.74 (Acc.)
16. Potthast et al. [potthast-etal-2018-stylometric] A Syntax, sentiment, readability (A.i) Buzzfeed-Webis Bag-of-words 0.46 (F1)
17. Reis et al. [reis_explainable_2019] A Syntax (A.i); contents (C.i); source rep. (U.i); timing (N.ii); URL (W.i) BuzzFeed GBM 0.85 (AUC)
18. Rubin et al. [rubin_towards_nodate] A Syntax (A.i) AMT Clustering 0.67 (Agreement)
19. Ruchansky et al. [ruchansky_csi_2017] A U (A.i); Responses (A.ii); acct metadata (U.i) Twitter/Weibo posts RNN/LSTM 0.95 (Acc.)
20. Santos et al. [santos_readability_2020] A Readability (A.i) Fake.Br corpus SVM 0.92 (Acc.)
21. Silva et al. [silva_embracing_2021] A N Topic detection (A.ii); propagation (N.i) PolitiFact[politifact], GossipCop[gossipcop], CoAID Clustering 0.88 (Acc.)
21. Singh et al. [singh_automated_nodate] A Syntax (A.i) Kaggle Fake News SVM 0.87 (Acc.)
22. Uppal et al. [uppal_fake_2020] A Discourse structure (A.i) Buzzfeed, PolitiFact[politifact] GRU, Dependency tree 0.74 (Acc.)
1. Cao et al. [cao_aiding_2012] U N Acct. cred. (U.i); prop. (N.i) Tuenti social network Louvain clustering 0.90+ (TP)
2. Danezis et al. [danezis2009sybilinfer] U N Acct. cred. (U.i); prop. (N.i) LiveJournal data Bayesian inf. n/a*
3. Ezzeddine et al. [ezzeddine2023exposing] U Acct. behaviors (U.ii) DATA LSTM 0.91 (AUC)
4. Hamdi et al. [hamdi_hybrid_2020] U N Account metadata (U.i); prop. (N.i) CREDBANK LDA, Bayes, Logit, SVM 0.99 (AUC)
5. Helmstetter et al. [helmstetter-supervised-2018] U N A Acct metadata (U.i); post sharing data (U.ii) Public site cred. lists SVM, NB, DT, RF 0.936 (F1)
6. Jain et al. [jain_towards_2016] U A Acct metadata (U.i); topic det. (A.ii) Twitter (API) Gensim, TextBlob 0.77 (Acc.)
7. Leonardi et al. [leonardi_users_2021] U N A Acct metadata (U.i); prop. (N.i) CoAID RF 0.81 (F1)
8. Saeed et al. [saeed2022trollmagnifier] U User behavior (U.ii) Reddit Pushshift; Reddit IRA trolls list RF 0.98 (Acc.)
9. Sansonetti et al. [sansonetti_unreliable_2020] U C Acct metadata (U.i); acct activity (U.ii) PolitiFact [politifact], Twitter (API) LSTM-CNN, SVM, KNN 0.92 (Acc.)
10. Santia et al. [santia_bots_2019] U A Source rep. (U.i); user response (U.ii); syntax (A.i) BuzzFeed SVM, RF, DT, NB 0.77 (Prec.)
11. Shu et al. [shu_exploiting_2017] U N A User behavior (U.ii); prop. (N.i) syntax (A.i) Buzzfeed, PolitiFact Gibbs sampling 0.85+ (Acc.)
12. Vargas et al. [vargas2020detection] N A Prop. (N.i); topic det. (A.ii) Twitter (API) RF 0.98 (F1)
13. Wang et al. [wang_sybil_2013] U N User behavior (U.ii) Renren data SVM 0.99 (Acc.)
14. Yu et al. [yu_sybillimit_2008] U N User behavior (U.i); prop. (N.i) LiveJournal, Friendster, DBLP accounts Random route n/a*
15. Yuan et al. [yuan2019detecting] U N Acct metadata (U.i); timing (N.ii) WeChat data Clustering 0.90+ (Prec.)
16. Zhang et al. [zhang_misinformation_2016] U N User behavior (U.ii); prop. (N.i) Twitter, Slashdot, Epinion Graph cut n/a
17. Zhou et al. [zhou_network-based_2019] U N User suscept. (U.i); prop (N.i) PolitiFact [politifact], BuzzFeed SVM, KNN, NB, DT, RF 0.93 (Acc.)
1. Alizadeh et al. [alizadeh2020content] N A U Propagation (N.i); syntax (A.i); acct metadata (U.i) Twitter (API), Reddit IRA troll list RF 0.70+ (F1)
2. Antoniadis et al. [antoniadis_model_2015] U A Acct metadata (U.i); syntax (A.i) Hurricane Sandy tweet dataset J48, RF, KNN, Bayes 0.79 (Avg. Prec.)
3. Assenmacher et al. [assenmacher2020two] N A Propagation (N.i); topic det. (A.ii) Twitter (API) Clustering not reported
4. Buntain et al. [buntain_automatically_2017] N U A Time (N.ii); acct metadata (U.i); sentiment (A.i) CREDBANK, Buzzfeed RF 0.65 (Acc.)
5. Castillo et al. [castillo_information_2011] U A Syntax (A.i); user behavior (U.ii) Twitter Monitor events SVM, DT 0.874 (P)
6. Chen et al. [chen_rumor_2018] N U A Social graph (N.iii); syntax (A.i); user behavior (U.ii) Weibo RNN 0.92 (Acc.)
7. Guo et al. [guo_rumor_2018] N C Prop. (N.i); semantics (C.i) Twitter, Weibo LSTM 0.9 (Acc.)
8. Jin et al. [jin_news_2016] N A Propagation (N.i); stance (A.iii) Sina Weibo posts Clustering 0.84 (Acc.)
9. Liu et al. [liu_early_nodate] N Propagation (N.i) Weibo, Twitter15, Twitter16 RNN, CNN 0.897 (Acc.)
10. Ma et al. [ma_detect_2017] N Propagation (N.i) Kochina, Ma, Shu Twitter datasets RNN/BiLSTM 0.75 (Acc.)
11. Magelinski et al. [magelinski_synch_2022] N U Prop. (N.i); timing (N.ii); user behavior (U.ii) Twitter (API) – n/a
12. Nguyen et al. [nguyen_fake_2019] N A Prop. (N.i); semantics (A.i) Twitter, Weibo, PHEME [pheme_2018] SVM, RNN, DT 0.970 (Acc.)
13. Pacheco et al. [pacheco_uncovering_2020] U N Account metadata (U.i); timing (N.i) Twitter (API) Clustering 0.8+ (Prec.)
14. Ratkiewicz et al. [ratkiewicz2011detecting] N A U Prop. (N.i); keywords (A.ii); user behavior (U.ii) Twitter (API) AdaBoost, SVM 0.96 (Acc.)
15. Sharma et al. [sharma_identifying_2021] N A U Prop. (N.i); keywords (A.ii); user behavior (U.ii) Twitter (API); Twitter IRA trolls list GMM, Kmeans, NN 0.94 (AUC)
16. Tschiatschek et al. [tschiatschek_fake_2018] N U Prop. (N.i); user rep. (U.i) Facebook dataset Bayes inf. n/a
17. Zeng et al. [zeng_unconfirmed_2016] N A Prop. (N.i); stance (A.iii) 4.3K tweets (from API) Logit, NB, RF 0.88 (Acc.)
1. Asr et al. [torabi_asr_data_2018] W A Source rep. (W.i); Syntax (A.i) BuzzfeedUSE, Snopes, Rashkin, Rubin CNN, SVM, NB not reported
2. Baly et al. [baly_predicting_nodate] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[mediabiasfactheck] SVM 0.66 (Acc.)
3. Baly et al. [baly-etal-2020-written] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[mediabiasfactheck] SVM 0.7152 (Acc.)
4. Castelo et al. [castelo_agnostic_2019] W A Site infra. (W.ii); syntax (A.i) Celebrity, US-Election2016 SVM, kNN, RF 0.86 (Acc.)
5. Chen et al. [chen_proactive_2020] W A Hosting infra. (URL) (W.ii); syntax (A.i) PoliticalFakeNews Clustering 0.97 (AUC)
6. Hounsel et al. [hounsel-infra-2020] W Site infra. (W.i) FactCheck, Snopes, PolitiFact, Buzzfeed RF 0.98 (AUC)
7. Ribeiro et al. [ribeiro_media_2018] W U Site infra. (W.i); User demog. (U.i) Facebook API Graph search 0.97 (PCC)
Table 4: Focus corpus by scope and target. Coding of a focus set of 87 papers, sorted by information scope. Values in parentheses in “Target” field correspond to highlighted subcategories presented in Section 4 (e.g., “C.i” denotes target (i) in “Claims,” Section 4.1). If authors present evaluation results for multiple models, we underline the most performant model and record its corresponding performance score.